# FactLens: Benchmarking Fine-Grained Fact Verification

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive capability in language generation and understanding, but their tendency to hallucinate and produce factually incorrect information remains a key limitation. To verify LLM-generated contents and claims from other sources, traditional verification approaches often rely on holistic models that assign a single factuality label to complex claims, potentially obscuring nuanced errors. In this paper, we advocate for a shift toward fine-grained verification, where complex claims are broken down into smaller sub-claims for individual verification, allowing for more precise identification of inaccuracies, improved transparency, and reduced ambiguity in evidence retrieval. However, generating sub-claims poses challenges, such as maintaining context and ensuring semantic equivalence with respect to the original claim. We introduce FactLens [^1] Link to our data and code, a benchmark for evaluating fine-grained fact verification, with metrics and automated evaluators of sub-claim quality. The benchmark data is manually curated to ensure high-quality ground truth. Our results show alignment between automated FactLens evaluators and human judgments, and we discuss the impact of sub-claim characteristics on the overall verification performance.
@@ 1 Introduction
Previous works on building fact-checking benchmarks focus on generating claims with a ground truth label, and in some cases provide the evidence/context to verify the claim. Aly et al. (2021); Schlichtkrull et al. (2023). Claims are generated using human annotators Aly et al. (2021), synthetic processes Bayat et al. (2023); Tang et al. (2024), or considering LLM outputs on Question-Answering tasks Wang et al. (2023). To increase the complexity of the fact-checking process, the claims are generated from source data of multiple domains & modalities, such as Wikipedia text and/or tables Thorne et al. (2018); Chen et al. (2019); Aly et al. (2021), Web Pages Schlichtkrull et al. (2023), Knowledge Graphs Kim et al. (2023), online posts/chats Wang et al. (2023); Li et al. (2023), and Q-A tasks from various domains such as statistics, finance, legal, etc Jacovi et al. (2024a).
@@
Despite this structured pipeline, most existing methods rely on a holistic verification model, where complex claims are assigned a single factuality label, often obscuring the nuanced nature of the errors or inaccuracies in the claims. In this work, we echo the sentiments of Wang et al. (2023); Liu et al. (2019); Pan et al. (2023) for a shift towards fine-grained verification of complex claims, where claims are decomposed into smaller, more manageable sub-claims that can be individually verified. We additionally emphasise on the need to provide evaluation metrics to benchmark such fine-grained verification and enrich existing benchmarks with fine-grained verification labels.
@@ 2.1 Evaluation Metrics
For each of these metrics, the sub claims are evaluated by assigning a score of ‘low’, ‘medium’ or ‘high’. For coverage and redundancy, the scores are assigned at the claim level as we consider the sub-claims as a whole. For all other metrics, scores are assigned to each sub-claim.
@@ 3.2 Evaluation Results
Figure 2: Impact of sub-claim quality metrics on verification. See Appendix E for details.
In fine-grained verification, our FactLens evaluators can act as judges of generated sub-claims, providing early revision signals if the sub-claims might lead to problematic verifications. To illustrate this, we perform an end-to-end evaluation (as shown in Figure 2) to highlight how the final verification accuracy is affected by the quality of the sub-claims. For example, we expect high-quality sub-claims to exhibit low fabrication scores. We note that for claim decompositions with a fabrication score classified as ‘low,’ the final fact-checking accuracy is higher compared to those with ‘medium’ or ‘high’ fabrication scores.
@@ B.2 Statistically Computed Evaluation
To compute the FactLens Evaluation metrics using statistical methods we rely on entity-based and semantic-based calculations. Given one instance, with claim $C$, we extract all the (Subject, Object) pairs present within it, and from there create a list of $S$ = subjects, and $O$ = objects. After decomposing the claim, we obtain sub-claims $c$ = {$c_{1}$, $c_{2}$, …, $c_{n}$}. For all $i$ in [1, n], we extract the subjects $s_{i}$ and object $o_{i}$ lists in a similar manner. We next follow these definitions to calculate the following metrics:
@@ B.3 FactLens Evaluation using Ensemble Method
B.4 Model Performance
Model Atomicity Sufficiency Fabrication Coverage Redundancy Readability
Llama-3.1 1.87 2.85 1.01 2.88 1.09 2.96
GPT-4o 1.82 2.85 1.02 2.89 1.15 2.95
Table 4: Measure of sub-claim quality using prompt-based contextualized decomposition. We report average scores on CoverBench dataset. We expect fabrication & redundancy to be low ($\sim$ 1), while other metrics to be high ($\sim$ 3).
We utilize the prompt defined in Table 2 to decompose claims using GPT-4o and Llama-3.1(405B). In Table 4 we tabulate the evaluation performance of both these models on the claim decomposition task using FactLens Evaluator. For each instance, we map the ‘low’, ‘medium’, ‘high’ scores (‘non-atomic-2’, ‘non-atomic-1’, ‘atomic’ for atomicity) to numerical values (1, 2, 3 respectively), and report the average for each metric across all 733 instances in the CoverBench dataset.
B.5 Agreement Scores on Synthetic Data

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive capability in language generation and understanding, but their tendency to hallucinate and produce factually incorrect information remains a key limitation. To verify LLM-generated contents and claims from other sources, traditional verification approaches often rely on holistic models that assign a single factuality label to complex claims, potentially obscuring nuanced errors. In this paper, we advocate for a shift toward fine-grained verification, where complex claims are broken down into smaller sub-claims for individual verification, allowing for more precise identification of inaccuracies, improved transparency, and reduced ambiguity in evidence retrieval. However, generating sub-claims poses challenges, such as maintaining context and ensuring semantic equivalence with respect to the original claim. We introduce FactLens111link to our data and code redacted for review , a benchmark for evaluating fine-grained fact verification, with metrics and automated evaluators of sub-claim quality. The benchmark data is manually curated to ensure high-quality ground truth. Our results show alignment between automated FactLens evaluators and human judgments, and we discuss the impact of sub-claim characteristics on the overall verification performance.
@@ 1 Introduction
Previous works on building fact-checking benchmarks focus on generating claims with a ground truth label, and in some cases provide the evidence/context to verify the claim, Aly et al. (2021); Schlichtkrull et al. (2023). Claims are generated using human annotators Aly et al. (2021), synthetic processes Bayat et al. (2023); Tang et al. (2024), or considering LLM outputs on Question-Answering tasks Wang et al. (2023). To increase the complexity of the fact-checking process, the claims are generated from source data of multiple domains & modalities, such as Wikipedia text and/or tables Thorne et al. (2018); Chen et al. (2019); Aly et al. (2021), Web Pages Schlichtkrull et al. (2023), Knowledge Graphs Kim et al. (2023), online posts/chats Wang et al. (2023); Li et al. (2023), and Q-A tasks from various domains such as statistics, finance, legal, etc Jacovi et al. (2024a).
@@
Despite this structured pipeline, most existing methods rely on a holistic verification model, where complex claims are assigned a single factuality label, often obscuring the nuanced nature of the errors or inaccuracies in the claims. In this work, we echo the sentiments of Wang et al. (2023); Liu et al. (2019); Pan et al. (2023); Min et al. (2023); Si et al. (2024) for a shift towards fine-grained verification of complex claims, where claims are decomposed into smaller, more manageable sub-claims that can be individually verified. We additionally emphasise on the need to provide evaluation metrics to benchmark such fine-grained verification and enrich existing benchmarks with fine-grained verification labels.
@@ 2.1 Evaluation Metrics
For each of these metrics, the sub claims are evaluated by assigning a score of ‘low’, ‘medium’ or ‘high’. For coverage and redundancy, the scores are assigned at the claim level as we consider the sub-claims as a whole. For all other metrics, scores are assigned to each sub-claim and then aggregated.
Ideal metric values: For an ideal claim decomposition, we expect the sub-claim to possess high atomicity, high sufficiency, high coverage and high readability while having low fabrication and low redundancy.
@@ 3.1 Dataset Creation
3.2 Claim Decomposition: Model Performance
Model Atomicity $\uparrow$ Sufficiency $\uparrow$ Fabrication $\downarrow$ Coverage $\uparrow$ Redundancy $\downarrow$ Readability $\uparrow$
Llama-3.1 1.87 2.85 1.01 2.88 1.09 2.96
GPT-4o 1.82 2.85 1.02 2.89 1.15 2.95
Table 2: Measure of sub-claim quality using prompt-based contextualized decomposition. We report average scores on CoverBench. Up & down arrows indicate which metrics should ideally be high ($\sim$ 3) or low ($\sim$ 1) respectively.
We utilize the prompt defined in Table 3 to decompose claims using GPT-4o and Llama-3.1(405B). In Table 2 we tabulate the evaluation performance of both these models on the claim decomposition task using FactLens Evaluator.
For each instance, we map the ‘low’, ‘medium’, ‘high’ scores (‘non-atomic-2’, ‘non-atomic-1’, ‘atomic’ for atomicity) to numerical values (1, 2, 3 respectively), and report the average for each metric across all 733 instances in the CoverBench dataset.
@@ 3.3 Evaluation Results
Figure 2: Impact of sub-claim quality metrics on verification performance. See Appendix E for details.
In fine-grained verification, our FactLens evaluators can act as judges of generated sub-claims, providing early revision signals if the sub-claims might lead to problematic verifications. To illustrate this, we perform an end-to-end evaluation (as shown in Figure 2) to highlight how the final verification performance is affected by the quality of the sub-claims. For example, we expect high-quality sub-claims to exhibit low fabrication scores. We note that for claim decompositions with a fabrication score classified as ‘low,’ the downstream fact-checking performance i.e. F1-score, is higher compared to those with ‘medium’ or ‘high’ fabrication scores.
@@ B.2 Statistically Computed Evaluation
To compute the FactLens Evaluation metrics using statistical methods we rely on entity-based and semantic-based calculations. Given one instance, with claim $C$, we extract all the (Subject, Object) pairs present within it using gpt-4o-mini; temperature = 0, and from there create a list of $S$ = subjects, and $O$ = objects. After decomposing the claim, we obtain sub-claims $c$ = {$c_{1}$, $c_{2}$, …, $c_{n}$}. For all $i$ in [1, n], we extract the subjects $s_{i}$ and object $o_{i}$ lists in a similar manner. We next follow these definitions to calculate the following metrics:
@@ B.3 FactLens Evaluation using Ensemble Method
B.4 Agreement Scores on Synthetic Data
In Table 1 we observed fair to moderate correlation between humans and FactLens Evaluator scores through Pearson and Spearman correlation scores. We provide the corresponding p-values in Table 5, from which we conclude the correlation scores for atomicity, coverage, fabrication (LLM-Human) and redundancy (LLM-Human) are statistically significant.
P-values atomicity sufficiency fabrication coverage redundancy
$r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$
LLM 4e-10 5e-10 0.1402 0.4367 5.1e-4 2.63e-7 2.9e-4 2.4e-4 3.26e-6 7.78e-6
Statistical 1e-9 1e-9 0.1656 0.1764 7.7e-8 1.1e-6 0.4043 0.3316
Table 5: P-values from Table 1